# Phase segregation lubricates paste extrusion

### v1

@@ A.4.1 Surface topology and wear suggest mixed regime lubrication
Surface topology scans of pellets were conducted, to support the hypothesis of decreased contact area between the two sliding surfaces as a result of a water film with increasing thickness between the pellet and metal surface. Providing evidence for the type of lubrication regime during pellet extrusion. At low water contents, distinct striations are visible along the direction of extrusion (Fig. A.6b&d), indicating a boundary regime lubrication with direct contact between the pellet and die-wall. Under these conditions, the calculated lubrication layer thickness is relatively small in comparison to the expected roughness of the metal die inside the holes (see Fig 1d), resulting in surface wear due to protrusion of the metal asperities into the softer pellets’ surface. We were unable to obtain a measure for the surface roughness of the die during these experiments. This would require cutting of the stainless steel die into measurable pieces and thus destroy valuable equipment. However, the striations disappear when increasing the water content from $0.009$ kg water per kg ingredients during the conditioning process to $0.057$ kg water per kg ingredients (Fig. A.6c&e). This suggests that, in this case, the lubrication layer thickness increases, filling more cavities in the asperities between the pellet and the metal die, transitioning into the mixed lubrication regime [25]. In the mixed regime, the contact between the two surfaces is reduced, resulting in less wear on the pellet’s surface.
Figure A.6: Increasing lubrication layer thickness reduces the real contact area between the pellet and the die, reducing surface wear. a, Schematic representation of the pellet sliding along the metal surface of the pelleting die with asperities. Increasing lubrication layers are expected to increase the separation between the pellet and the die’s surface, reducing their contact area as expected in the mixed lubrication regime [25]. b & c, 3D surface topology scans of two pellets produced at low $\Delta\mathrm{mc}$ and high $\Delta\mathrm{mc}$, $0.009$ and $0.057$ kg water per kg ingredients respectively. d & e are the surface height profiles of the pellets scanned in (b) and (c). The black area’s correspond to peaks and valleys with a height or depth larger than $1.5\mathrm{\mu m}$. In (b) and (d) striations along the sliding direction are clearly visible, due to a thin lubrication layer. In (c) and (e) those striations are absent, due to filling of the cavities inside the pellet and the die, resulting in a reduced protrusion of the metal asperities into the extruded pellet.
@@ A.4.2 Mixed lubrication and the apparent coefficient of friction
Here, $X$ (-), with values from 00 and $1$, represents the ratio between the load carried by the solid-solid contact interface with respect to the total contact load. The remainder, $(1-X)$, represents the load carried by the fluid film between the two surfaces. In the case of boundary lubrication, $X$ is equal to 1, and the apparent coefficient of friction, $\mu$, is fully determined by the solid-solid sliding friction. Non-lubricated pellet extrusion would occur during pellet production without the addition of water, steam, fat, oil, or other fluid lubricants. However, in the presence of a water film between the two sliding surfaces, $X$ reduces to values smaller than $1$, and the apparent coefficient of friction, determined from the experimental processing data, decreases.
A.4.3 Experimental coefficient of friction and energy use
According to the work by Holm et al., published in [45] and [47], a theoretical relationship exists that couples compression ratio - die length over die diameter, $L/D$ (-), and the coefficient of friction, $\mathrm{\mu}$, to the pressure required for extrusion during pelletization. In their works Holm et al. derive an exponential relationship (see eq. 3)
$P_{x}(x)=\frac{P_{0}}{\nu}\ \exp(4\mu\nu\frac{L}{D}-1)$ (3)
In which the extrusion pressure, $P_{x}$ $(\mathrm{Pa})$, is determined by a prestressing constant, $P_{0}$ $(\mathrm{Pa})$, the radial-to-axial stress ratio constant, or lateral pressure coefficient , $\nu$ (-), the coefficient of friction and the compression ratio of the die (see refs [45, 63] for further reference). By taking $\nu$ as a ingredient specific constant during pelletization, the pressure for extrusion increases exponentially with $\mu$ and $L/D$. Furthermore the extrusion pressure $(\mathrm{Nm^{-2}})$ and power consumption $(\mathrm{N\ ms^{-1}=W})$ are related through the volume flow rate $(\mathrm{m^{3}s^{-1}})$. Consequently, the power consumption of the pellet press, as well as the specific mechanical energy usage - the energy per unit mass - is expected to increase exponentially as a function of $\mu$ and $L/D$ as well (Fig. A.7).
Figure A.7: Increasing steam addition reduces the coefficient of friction between the pellet and the metal die. a, During pellet extrusion the specific mechanical energy usage is exponentially proportional to the coefficient of friction and compression ratio of the pelleting die. b Increasing the steam use during steam conditioning from a low moisture content ($0.035$ kg steam per kg ingredients) to a higher moisture content ($0.053$ kg steam per kg ingredients), results in a lower energy usage due to a decrease in the amount of friction. This is consistent with previous observations [37, 1].
Increasing the steam use during steam conditioning, achieved by controlling the temperature of the ingredient mixture after conditioning, leads to higher moisture content in the ingredient mixture and, as a result, reduces the mechanical energy demand during pellet extrusion [37, 1]. Our experimental data indicates that increasing steam use from $0.035$ kg steam per kg ingredients to $0.053$ kg steam per kg ingredients results in a reduction of the apparent coefficient of friction by approximately $16\%$ (see Fig. A.7). The thickness of the lubricating film is directly related to the increased separation of the two sliding surfaces (Fig 1d&e). By increasing the lubrication layer thickness the direct contact area between the metal die and the sliding pellet reduces, consequently $X$ reduces (see previous section) and the relative contribution of $\mu_{\mathrm{solid}}$ with respect to the apparent friction coefficient, $\mu$, decreases, reducing the extrusion sliding resistance. The contribution of the friction within the water film, is negligible (appendix C.2 - Energy dissipation in a liquid film). Hence our experimental observations supports the hypothesis that the increasing use of steam, and consequently the increase in moisture content, contributes to lowering the coefficient of friction between the metal die and the sliding pellet, through the formation of a lubrication layer in the mixed lubrication regime.
A.4.4 The effect of production rate on the mechanical energy use
@@
Figure A.8: Mechanical energy use during pellet extrusion decreases with production rate in the mixed lubrication regime. The transition into the mixed lubrication regime is marked by a reduction in mechanical energy use as the sliding velocity (or production rate) increases, corresponding to a decrease in the coefficient of friction. The error-bars correspond to the minimum and maximum values of two or three independent weighing samples taken during experimental runs.
@@ A.5 Physical pellet quality
Figure A.9: Physical pellet quality can be optimized by tuning pellet temperature through steam and friction. a, During pelleting, energy is required for the production of steam (marked by “o”). In addition, electrical motor power is dissipated into the pellets (marked by “+”). Here, the total mechanical energy dissipated by the press, including the idle load of the pellet press (gross mechanical energy), is represented. The sum of the steam energy and mechanical energy determines the total energy used in pellet production (marked by “•”), which includes both conditioning and pelletizing. An optimum energy use exists, coupled with a normative pellet durability value (here $0.9$, equal to $90\%$), specifying the minimum amount of energy required to produce a high-quality pellet. This optimum is reached when pellet durability (b) no longer increases significantly with the invested energy during production (a). Below this optimum, pellets are still formed; however, the quality is insufficient for downstream processes, while above this optimum, the quality increases disproportionately with the energy investment (b). This pellet durability increase with increasing steam usage is consistent with [37].
@@
Producing a durable pellet with minimal energy use involves precise control over heat fluxes within the pellet production line. This control allows efficient adjustment of ingredient and pellet temperatures to facilitate bond formation and produce durable pellets (see Fig. A.9b). The heating process begins with the addition of steam during steam conditioning, followed by heating through friction during pellet extrusion. Increasing steam usage enables a reduction in mechanical energy (see Fig. 3b and Fig. A.9a). However, it is essential to note that steam production itself is energy-intensive (see Fig. A.9a), here estimated through assuming a $100\%$ efficiency of the steam-boiler. Therefore, finding a balance between steam and mechanical energy fluxes is crucial for optimizing pellet production for energy efficiency. Beyond a certain temperature threshold for a specified ingredient composition, further increases in total energy input (see Fig. A.9a - ”Total energy use”) yield no significant durability improvements (see Fig. A.9b). Above this optimum process temperature, production costs increase significantly with continued steam usage. To mitigate these costs, replacing part of the added steam with tap water during conditioning might be considered. Tap water can contribute to lubrication, reducing mechanical energy use without the energy costs associated with an increased steam production. Therefore active monitoring of the lubrication layer, the mechanical energy use and product temperatures, is a requirement for future process optimization during pellet extrusion.
Appendix B Methods
@@ B.1 Materials
A total of 18,000 kg of milled material was obtained from Research Diet Services (Wijk bij Duurstede, The Netherlands). The material was ground using a hammer mill equipped with 4-mm screens. The material, a 50/50% w/w mixture of cleaned maize kernels and low-sugar sugar beet pulp, was delivered in two separate orders: the first batch (Batch 1) contained 7,000 kg, and the second batch (Batch 2) contained 11,000 kg. Batch 1 was used to obtain the results presented in all figures except Fig. A.7, Batch 2 was used solely for the results presented in Fig. A.7. The products, being of organic/natural origin, may differ in their chemical compositional between Batch 1 and Batch 2. However, these variations were not the focus of this study, therefore the chemical composition was not analyzed.
@@ B.2 Pellet production and process monitoring
A RMP200 ring-die pelletizer (Münch Edelstahl GmbH, Hilden Germany) served as the pilot-scale pellet extrusion system. Configured with two rollers, the ring-die featured 6 mm diameter die-holes and a compression ratio (L/D) of 12, corresponding to a channel length of 72 mm. In addition to the 6 mm L/D 12 die, three additional dies with 6 mm diameter holes were used for the results presented in Fig. A.7, with L/D ratio’s 6, 9.3 and 16 respectively. The roller-die gap was set at 0.20 mm, and the die rotation frequency maintained at 5 Hz.
@@
During the processing of Batch 1, the effect of steam dosing on pellet extrusion was evaluated by conducting five steady-state runs in random order at constant production capacity (approx. 250 kg/h) and variable steam injection rate, thereby tuning $\Delta\mathrm{mc}$ as specified in Fig. 1d. During the processing of Batch 2, the effect of L/D ratio and steam dosage on the specific mechanical energy use was evaluated using a constant production capacity (approx. 250 kg/h) and at two levels of $\Delta\mathrm{mc}$ (0.035 and 0.053 kg steam per kg ingredients, respectively), while changing the die-configuration as specified in Fig. A.7b.
@@ C.1 Simulation details COMSOL
All simulations were initialized with an initial moisture content of $+2.5\%$ (kg water per kg material), a typical increase in moisture content during steam conditioning.
@@ C.2 Energy dissipation in a liquid film
The surface topology scans, exhibiting clear striation patterns upon the addition of minimal water during pellet manufacturing and the disappearance thereof with increased water addition, suggest that we are within the mixed lubrication regime. In this regime, where $X$ ranges between $1$ and 00, the work of friction is predominantly determined by solid-solid contact friction (see eq. 2). The exact contributions of $\mu_{\mathrm{solid}}$ and $\mu_{\mathrm{water}}$ at operational processes are unknown. However, the contribution of friction in the water film can be estimated by assuming a full separation of the two sliding surfaces, in which case all friction is assumed to result from shearing of the lubrication layer. Through careful estimates, we can determine the specific energy dissipation due to the shearing of the water film (see eq. 7):

### v2

@@ A.4.1 Surface topology and wear suggest mixed regime lubrication
Surface topology scans of pellets were conducted, to support the hypothesis of decreased contact area between the two sliding surfaces as a result of a water film with increasing thickness between the pellet and metal surface. Providing evidence for the type of lubrication regime during pellet extrusion. At low water contents, distinct striations are visible along the direction of extrusion (Fig. A.6b&d), indicating a boundary regime lubrication with direct contact between the pellet and die-wall. Under these conditions, the calculated lubrication layer thickness is relatively small in comparison to the expected roughness of the metal die inside the holes (see Fig 2d), resulting in surface wear due to protrusion of the metal asperities into the softer pellets’ surface. We were unable to obtain a measure for the surface roughness of the die during these experiments. This would require cutting of the stainless steel die into measurable pieces and thus destroy valuable equipment. However, the striations disappear when increasing the water content from 0.0090.0090.009 kg water per kg ingredients during the conditioning process to 0.0570.0570.057 kg water per kg ingredients (Fig. A.6c&e). This suggests that, in this case, the lubrication layer thickness increases, filling more cavities in the asperities between the pellet and the metal die, transitioning into the mixed lubrication regime [28]. In the mixed regime, the contact between the two surfaces is reduced, resulting in less wear on the pellet’s surface.
Figure A.6: (a) Schematic representation of the pellet sliding along the metal surface of the pelleting die with asperities. Increasing lubrication layers are expected to increase the separation between the pellet and the die’s surface, reducing their contact area as expected in the mixed lubrication regime [28]. (b,c) 3D surface topology scans of two pellets produced at low $\Delta\mathrm{mc}$ and high $\Delta\mathrm{mc}$, 0.0090.0090.009 and 0.0570.0570.057 kg water per kg ingredients respectively. (d) & (e) are the surface height profiles of the pellets scanned in (b) and (c). The black area’s correspond to peaks and valleys with a height or depth larger than $1.5\mathrm{\mu m}$. In (b) and (d) striations along the sliding direction are clearly visible, due to a thin lubrication layer. In (c) and (e) those striations are absent, due to filling of the cavities inside the pellet and the die, resulting in a reduced protrusion of the metal asperities into the extruded pellet.
@@ A.4.2 Mixed lubrication and the apparent coefficient of friction
Here, $X$ (-), with values from 0 and 111, represents the ratio between the load carried by the solid-solid contact interface with respect to the total contact load. The remainder, $(1-X)$, represents the load carried by the fluid film between the two surfaces. In the case of boundary lubrication, $X$ is equal to 1, and the apparent coefficient of friction, $\mu$, is fully determined by the solid-solid sliding friction. Non-lubricated pellet extrusion would occur during pellet production without the addition of water, steam, fat, oil, or other fluid lubricants. However, in the presence of a water film between the two sliding surfaces, $X$ reduces to values smaller than 111, and the apparent coefficient of friction, determined from the experimental processing data, decreases.
A.4.3 The effect of production rate on the mechanical energy use
@@
Figure A.7: The transition into the mixed lubrication regime is marked by a reduction in mechanical energy use ($SME_{\rm net}$) as the sliding velocity (or production rate) increases, corresponding to a decrease in the coefficient of friction, as observed. The error-bars correspond to the minimum and maximum values of two or three independent weighing samples taken during experimental runs.
@@ A.5 Physical pellet quality
[…]
@@
[…]
Appendix B Methods & Protocols
@@ B.1 Materials
A total of 18,000 kg of milled material was obtained from Research Diet Services (Wijk bij Duurstede, The Netherlands). The material was ground using a hammer mill equipped with 4-mm screens. The material, a 50/50% w/w mixture of cleaned maize kernels and low-sugar sugar beet pulp, was delivered in two separate orders: the first batch (Batch 1) contained 7,000 kg, and the second batch (Batch 2) contained 11,000 kg. Batch 1 was used to obtain the results presented in all figures except Fig. 3, Batch 2 was used solely for the results presented in Fig. 3. The products, being of organic/natural origin, may differ in their chemical compositional between Batch 1 and Batch 2. However, these variations were not the focus of this study, therefore the chemical composition was not analyzed.
@@ B.2 Pellet production and process monitoring
A RMP200 ring-die pelletizer (Münch Edelstahl GmbH, Hilden Germany) served as the pilot-scale pellet extrusion system. Configured with two rollers, the ring-die featured 6 mm diameter die-holes and a compression ratio (L/D) of 12, corresponding to a channel length of 72 mm. In addition to the 6 mm L/D 12 die, three additional dies with 6 mm diameter holes were used for the results presented in Fig. 3, with L/D ratio’s 6, 9.3 and 16 respectively. The roller-die gap was set at 0.20 mm, and the die rotation frequency maintained at 5 Hz.
@@
During the processing of Batch 1, the effect of steam dosing on pellet extrusion was evaluated by conducting five steady-state runs in random order at constant production capacity (approx. 250 kg/h) and variable steam injection rate, thereby tuning $\Delta\mathrm{mc}$ as specified in Fig. 2d. During the processing of Batch 2, the effect of L/D ratio and steam dosage on the specific mechanical energy use was evaluated using a constant production capacity (approx. 250 kg/h) and at two levels of $\Delta\mathrm{mc}$ (0.035 and 0.053 kg steam per kg ingredients, respectively), while changing the die-configuration as specified in Fig. 3b.
@@ C.1 Simulation details COMSOL
[…]
@@ C.2 Energy dissipation in a liquid film
The surface topology scans, exhibiting clear striation patterns upon the addition of minimal water during pellet manufacturing and the disappearance thereof with increased water addition, suggest that we are within the mixed lubrication regime. In this regime, where $X$ ranges between 111 and 0, the work of friction is predominantly determined by solid-solid contact friction (see eq. 3). The exact contributions of $\mu_{\mathrm{solid}}$ and $\mu_{\mathrm{water}}$ at operational processes are unknown. However, the contribution of friction in the water film can be estimated by assuming a full separation of the two sliding surfaces, in which case all friction is assumed to result from shearing of the lubrication layer. Through careful estimates, we can determine the specific energy dissipation due to the shearing of the water film (see eq. 7):